# Generalize across Homophily and Heterophily: Hybrid Spectral Graph Pre-Training and Prompt Tuning

### v2

@@ E.4 Additional Accuracy Results
GCN
0.2535
$\pm$0.0611
0.1835
$\pm$0.0302
0.4028
$\pm$0.1002
0.2506
$\pm$0.0571
0.3782
$\pm$0.0853
0.2496
$\pm$0.0411
0.3117
$\pm$0.0532
0.2998
$\pm$0.0510
0.2541
$\pm$0.0123
0.2441
$\pm$0.0141
GAT
0.2620
$\pm$0.1373
0.1531
$\pm$0.0747
0.2986
$\pm$0.0242
0.2144
$\pm$0.0471
0.2891
$\pm$0.0961
0.1836
$\pm$0.0463
0.2989
$\pm$0.0489
0.2873
$\pm$0.0462
0.2307
$\pm$0.0067
0.2281
$\pm$0.0016
BernNet
0.3042
$\pm$0.0727
0.2287
$\pm$0.0393
0.5915
$\pm$0.0309
0.3743
$\pm$0.0476
0.2950
$\pm$0.0536
0.2215
$\pm$0.0343
0.2785
$\pm$0.0134
0.2745
$\pm$0.0139
0.2161
$\pm$0.0112
0.2150
$\pm$0.0108
ChebNetII
0.3296
$\pm$0.0433
0.2080
$\pm$0.0288
0.5606
$\pm$0.0559
0.3558
$\pm$0.0487
0.3584
$\pm$0.0462
0.2533
$\pm$0.0525
0.2718
$\pm$0.0247
0.2684
$\pm$0.0265
0.2144
$\pm$0.0123
0.2133
$\pm$0.0123
TFE-GNN
0.4535
$\pm$0.1001
0.3950 $\pm$0.1131
0.4563
$\pm$0.1230
0.3730
$\pm$0.1375
0.4733
$\pm$0.0973
0.3828
$\pm$0.0822
0.2907
$\pm$0.0091
0.2872
$\pm$0.0115
0.2256
$\pm$0.0056
0.2242
$\pm$0.0050
DGI
0.2648
$\pm$0.0552
0.1650
$\pm$0.0311
0.4789
$\pm$0.1321
0.3031
$\pm$0.0804
0.3545
$\pm$0.0967
0.2293
$\pm$0.0422
0.2907
$\pm$0.0388
0.2692
$\pm$0.0336
0.2274
$\pm$0.0301
0.1989
$\pm$0.0195
GraphCL
0.2648
$\pm$0.1258
0.1826
$\pm$0.0579
0.4676
$\pm$0.0833
0.2673
$\pm$0.0522
0.3386
$\pm$0.0731
0.2414
$\pm$0.0313
0.2921
$\pm$0.0475
0.2710
$\pm$0.0432
0.2215
$\pm$0.0066
0.1890
$\pm$0.0129
SimGRACE
0.3296
$\pm$0.0738
0.1904
$\pm$0.0237
0.4732
$\pm$0.0423
0.2833
$\pm$0.0187
0.3129
$\pm$0.0926
0.2116
$\pm$0.0625
0.3022
$\pm$0.0275
0.2706
$\pm$0.0228
0.2261
$\pm$0.0164
0.2071
$\pm$0.0165
PolyGCL
0.3070
$\pm$0.0559
0.2268
$\pm$0.0469
0.6338
$\pm$0.0570
0.4913
$\pm$0.0568
0.3644
$\pm$0.0820
0.2254
$\pm$0.0273
0.3415 $\pm$0.0379
0.3308
$\pm$0.0351
0.2516
$\pm$0.0170
0.2450
$\pm$0.0129
GPPT
0.3352
$\pm$0.0491
0.1798
$\pm$0.0447
0.5634
$\pm$0.0527
0.3247
$\pm$0.0407
0.3426
$\pm$0.0971
0.2619
$\pm$0.0726
0.3092
$\pm$0.0191
0.2976
$\pm$0.0221
0.2163
$\pm$0.0142
0.2082
$\pm$0.0187
GPrompt
0.3437
$\pm$0.1262
0.1918
$\pm$0.0365
0.3634
$\pm$0.2001
0.1932
$\pm$0.0846
0.3505
$\pm$0.0860
0.1989
$\pm$0.0425
0.2539
$\pm$0.0159
0.2326
$\pm$0.0144
0.2088
$\pm$0.0197
0.2019
$\pm$0.0176
GPF-plus
0.3549
$\pm$0.1629
0.1956
$\pm$0.0716
0.4761
$\pm$0.0449
0.2515
$\pm$0.0353
0.2891
$\pm$0.1204
0.1928
$\pm$0.0688
0.2467
$\pm$0.0246
0.2317
$\pm$0.0245
0.2076
$\pm$0.0121
0.1844
$\pm$0.0187
All-in-One
0.3662
$\pm$0.1718
0.1479
$\pm$0.0554
0.4028
$\pm$0.0250
0.1825
$\pm$0.0650
0.3069
$\pm$0.1693
0.1381
$\pm$0.0731
0.2417
$\pm$0.0180
0.2254
$\pm$0.0164
0.2249
$\pm$0.0530
0.1751
$\pm$0.0324
ProNoG
0.2563
$\pm$0.0449
0.1987
$\pm$0.0289
0.5070
$\pm$0.0807
0.2627
$\pm$0.0420
0.3366
$\pm$0.0349
0.2218
$\pm$0.0371
0.2803
$\pm$0.0296
0.2565
$\pm$0.0294
0.2118
$\pm$0.0072
0.2030
$\pm$0.0049
HS-GPPT 0.5239 $\pm$0.0675
0.3875
$\pm$0.0777
0.6423 $\pm$0.1041 0.5724 $\pm$0.0446 0.5683 $\pm$0.0640 0.4554 $\pm$0.0484
0.3302
$\pm$0.0357
0.3324 $\pm$0.0360 0.2589 $\pm$0.0144 0.2536 $\pm$0.0108
@@
DGI
0.3485
$\pm$0.1185
0.2123
$\pm$0.0627
0.3408
$\pm$0.1059
0.1976
$\pm$0.0721
0.2240
$\pm$0.0142
0.2048
$\pm$0.0093
0.4423
$\pm$0.1526
0.2177
$\pm$0.0762
0.2423
$\pm$0.0402
0.1450
$\pm$0.0258
0.1008
$\pm$0.0204
0.0615
$\pm$0.0075
GraphCL
0.1941
$\pm$0.0495
0.1433
$\pm$0.0313
0.5127
$\pm$0.1274
0.2847
$\pm$0.0680
0.2245
$\pm$0.0198
0.2098
$\pm$0.0164
0.4366
$\pm$0.0782
0.2684
$\pm$0.0266
0.2282
$\pm$0.0620
0.1686
$\pm$0.0438
0.0990
$\pm$0.0199
0.0681
$\pm$0.0090
SimGRACE
0.3228
$\pm$0.0774
0.1957
$\pm$0.0294
0.3465
$\pm$0.1293
0.2309
$\pm$0.0499
0.2276
$\pm$0.0140
0.2111
$\pm$0.0141
0.3408
$\pm$0.1074
0.1834
$\pm$0.0438
0.1972
$\pm$0.0891
0.1369
$\pm$0.0720
0.0958
$\pm$0.0186
0.0753
$\pm$0.0098
PolyGCL
0.3960
$\pm$0.0689
0.2473
$\pm$0.0460
0.5549
$\pm$0.0423
0.3901
$\pm$0.0585
0.2218
$\pm$0.0213
0.2130
$\pm$0.0217
0.4366
$\pm$0.1061
0.2934
$\pm$0.0774
0.3324
$\pm$0.1115
0.2112
$\pm$0.0410
0.0593
$\pm$0.0043
0.0487
$\pm$0.0020
GPPT
0.3683
$\pm$0.0637
0.2295
$\pm$0.0564
0.5437
$\pm$0.0211
0.3774
$\pm$0.0678
0.2155
$\pm$0.0167
0.1987
$\pm$0.0189
0.5437
$\pm$0.0211
0.3774
$\pm$0.0678
0.2282
$\pm$0.0522
0.1906
$\pm$0.0435
0.0990
$\pm$0.0155
0.0860
$\pm$0.0049
GPrompt
0.2911
$\pm$0.1204
0.1960
$\pm$0.0614
0.3662
$\pm$0.1336
0.2078
$\pm$0.0608
0.2042
$\pm$0.0125
0.1982
$\pm$0.0091
0.3972
$\pm$0.1467
0.2131
$\pm$0.0386
0.3408
$\pm$0.1139
0.1825
$\pm$0.0435
0.0641
$\pm$0.0032
0.0524
$\pm$0.0022
GPF-plus
0.3109
$\pm$0.1086
0.2046
$\pm$0.0736
0.5239
$\pm$0.0720
0.3112
$\pm$0.0626
0.2112
$\pm$0.0102
0.1808
$\pm$0.0216
0.5042
$\pm$0.0559
0.3323
$\pm$0.0347
0.3606
$\pm$0.1570
0.2559
$\pm$0.0941
0.0687
$\pm$0.0078
0.0519
$\pm$0.0031
All-in-One
0.2337
$\pm$0.0878
0.0982
$\pm$0.0418
0.4310
$\pm$0.2792
0.2317
$\pm$0.1106
0.2116
$\pm$0.0217
0.1902
$\pm$0.0203
0.3127
$\pm$0.2199
0.1616
$\pm$0.1077
0.3831
$\pm$0.1751
0.1829
$\pm$0.0666
0.0559
$\pm$0.0251
0.0152
$\pm$0.0057
ProNoG
0.3366
$\pm$0.0349
0.2218
$\pm$0.0371
0.2789
$\pm$0.0466
0.2060
$\pm$0.0290
0.2154
$\pm$0.0083
0.2064
$\pm$0.0113
0.2620
$\pm$0.1002
0.2155
$\pm$0.0666
0.2986
$\pm$0.1067
0.1869
$\pm$0.0382
0.0748
$\pm$0.0159
0.0566
$\pm$0.0144
@@
E.5 Results of Baselines with Other Backbones
@@
In this section, we present the results of the baselines that use a GraphTransformer model UniMP (Shi et al. 2020) and our used BWGNN (Tang et al. 2022) as their backbone. To avoid excessive computational costs from grid searches, we utilize optimal pre-training and prompt-tuning combinations from Tables 9, Table 10, and Table 11. The results of transductive learning are shown in Table 12 and 14, and those of inductive learning are presented in Table 13 and 15. The results indicate that our model outperforms the baselines with various backbones.
E.6 Ablation Study

### v3

@@ E.3 Additional Accuracy Results
GCN 0.2535 $\pm$0.0611 0.1835 $\pm$0.0302 0.4028 $\pm$0.1002 0.2506 $\pm$0.0571 0.3782 $\pm$0.0853 0.2496 $\pm$0.0411 0.3117 $\pm$0.0532 0.2998 $\pm$0.0510 0.2541 $\pm$0.0123 0.2441 $\pm$0.0141
GAT 0.2620 $\pm$0.1373 0.1531 $\pm$0.0747 0.2986 $\pm$0.0242 0.2144 $\pm$0.0471 0.2891 $\pm$0.0961 0.1836 $\pm$0.0463 0.2989 $\pm$0.0489 0.2873 $\pm$0.0462 0.2307 $\pm$0.0067 0.2281 $\pm$0.0016
BernNet 0.3042 $\pm$0.0727 0.2287 $\pm$0.0393 0.5915 $\pm$0.0309 0.3743 $\pm$0.0476 0.2950 $\pm$0.0536 0.2215 $\pm$0.0343 0.2785 $\pm$0.0134 0.2745 $\pm$0.0139 0.2161 $\pm$0.0112 0.2150 $\pm$0.0108
ChebNetII 0.3296 $\pm$0.0433 0.2080 $\pm$0.0288 0.5606 $\pm$0.0559 0.3558 $\pm$0.0487 0.3584 $\pm$0.0462 0.2533 $\pm$0.0525 0.2718 $\pm$0.0247 0.2684 $\pm$0.0265 0.2144 $\pm$0.0123 0.2133 $\pm$0.0123
TFE-GNN 0.4535 $\pm$0.1001 0.3950 $\pm$0.1131 0.4563 $\pm$0.1230 0.3730 $\pm$0.1375 0.4733 $\pm$0.0973 0.3828 $\pm$0.0822 0.2907 $\pm$0.0091 0.2872 $\pm$0.0115 0.2256 $\pm$0.0056 0.2242 $\pm$0.0050
DGI 0.2648 $\pm$0.0552 0.1650 $\pm$0.0311 0.4789 $\pm$0.1321 0.3031 $\pm$0.0804 0.3545 $\pm$0.0967 0.2293 $\pm$0.0422 0.2907 $\pm$0.0388 0.2692 $\pm$0.0336 0.2274 $\pm$0.0301 0.1989 $\pm$0.0195
GraphCL 0.2648 $\pm$0.1258 0.1826 $\pm$0.0579 0.4676 $\pm$0.0833 0.2673 $\pm$0.0522 0.3386 $\pm$0.0731 0.2414 $\pm$0.0313 0.2921 $\pm$0.0475 0.2710 $\pm$0.0432 0.2215 $\pm$0.0066 0.1890 $\pm$0.0129
SimGRACE 0.3296 $\pm$0.0738 0.1904 $\pm$0.0237 0.4732 $\pm$0.0423 0.2833 $\pm$0.0187 0.3129 $\pm$0.0926 0.2116 $\pm$0.0625 0.3022 $\pm$0.0275 0.2706 $\pm$0.0228 0.2261 $\pm$0.0164 0.2071 $\pm$0.0165
PolyGCL 0.3070 $\pm$0.0559 0.2268 $\pm$0.0469 0.6338 $\pm$0.0570 0.4913 $\pm$0.0568 0.3644 $\pm$0.0820 0.2254 $\pm$0.0273 0.3415 $\pm$0.0379 0.3308 $\pm$0.0351 0.2516 $\pm$0.0170 0.2450 $\pm$0.0129
GPPT 0.3352 $\pm$0.0491 0.1798 $\pm$0.0447 0.5634 $\pm$0.0527 0.3247 $\pm$0.0407 0.3426 $\pm$0.0971 0.2619 $\pm$0.0726 0.3092 $\pm$0.0191 0.2976 $\pm$0.0221 0.2163 $\pm$0.0142 0.2082 $\pm$0.0187
GPrompt 0.3437 $\pm$0.1262 0.1918 $\pm$0.0365 0.3634 $\pm$0.2001 0.1932 $\pm$0.0846 0.3505 $\pm$0.0860 0.1989 $\pm$0.0425 0.2539 $\pm$0.0159 0.2326 $\pm$0.0144 0.2088 $\pm$0.0197 0.2019 $\pm$0.0176
GPF-plus 0.3549 $\pm$0.1629 0.1956 $\pm$0.0716 0.4761 $\pm$0.0449 0.2515 $\pm$0.0353 0.2891 $\pm$0.1204 0.1928 $\pm$0.0688 0.2467 $\pm$0.0246 0.2317 $\pm$0.0245 0.2076 $\pm$0.0121 0.1844 $\pm$0.0187
All-in-One 0.3662 $\pm$0.1718 0.1479 $\pm$0.0554 0.4028 $\pm$0.0250 0.1825 $\pm$0.0650 0.3069 $\pm$0.1693 0.1381 $\pm$0.0731 0.2417 $\pm$0.0180 0.2254 $\pm$0.0164 0.2249 $\pm$0.0530 0.1751 $\pm$0.0324
ProNoG 0.2563 $\pm$0.0449 0.1987 $\pm$0.0289 0.5070 $\pm$0.0807 0.2627 $\pm$0.0420 0.3366 $\pm$0.0349 0.2218 $\pm$0.0371 0.2803 $\pm$0.0296 0.2565 $\pm$0.0294 0.2118 $\pm$0.0072 0.2030 $\pm$0.0049
HS-GPPT 0.5239 $\pm$0.0675 0.3875 $\pm$0.0777 0.6423 $\pm$0.1041 0.5724 $\pm$0.0446 0.5683 $\pm$0.0640 0.4554 $\pm$0.0484 0.3302 $\pm$0.0357 0.3324 $\pm$0.0360 0.2589 $\pm$0.0144 0.2536 $\pm$0.0108
@@
DGI 0.3485 $\pm$0.1185 0.2123 $\pm$0.0627 0.3408 $\pm$0.1059 0.1976 $\pm$0.0721 0.2240 $\pm$0.0142 0.2048 $\pm$0.0093 0.4423 $\pm$0.1526 0.2177 $\pm$0.0762 0.2423 $\pm$0.0402 0.1450 $\pm$0.0258 0.1008 $\pm$0.0204 0.0615 $\pm$0.0075
GraphCL 0.1941 $\pm$0.0495 0.1433 $\pm$0.0313 0.5127 $\pm$0.1274 0.2847 $\pm$0.0680 0.2245 $\pm$0.0198 0.2098 $\pm$0.0164 0.4366 $\pm$0.0782 0.2684 $\pm$0.0266 0.2282 $\pm$0.0620 0.1686 $\pm$0.0438 0.0990 $\pm$0.0199 0.0681 $\pm$0.0090
SimGRACE 0.3228 $\pm$0.0774 0.1957 $\pm$0.0294 0.3465 $\pm$0.1293 0.2309 $\pm$0.0499 0.2276 $\pm$0.0140 0.2111 $\pm$0.0141 0.3408 $\pm$0.1074 0.1834 $\pm$0.0438 0.1972 $\pm$0.0891 0.1369 $\pm$0.0720 0.0958 $\pm$0.0186 0.0753 $\pm$0.0098
PolyGCL 0.3960 $\pm$0.0689 0.2473 $\pm$0.0460 0.5549 $\pm$0.0423 0.3901 $\pm$0.0585 0.2218 $\pm$0.0213 0.2130 $\pm$0.0217 0.4366 $\pm$0.1061 0.2934 $\pm$0.0774 0.3324 $\pm$0.1115 0.2112 $\pm$0.0410 0.0593 $\pm$0.0043 0.0487 $\pm$0.0020
GPPT 0.3683 $\pm$0.0637 0.2295 $\pm$0.0564 0.5437 $\pm$0.0211 0.3774 $\pm$0.0678 0.2155 $\pm$0.0167 0.1987 $\pm$0.0189 0.5437 $\pm$0.0211 0.3774 $\pm$0.0678 0.2282 $\pm$0.0522 0.1906 $\pm$0.0435 0.0990 $\pm$0.0155 0.0860 $\pm$0.0049
GPrompt 0.2911 $\pm$0.1204 0.1960 $\pm$0.0614 0.3662 $\pm$0.1336 0.2078 $\pm$0.0608 0.2042 $\pm$0.0125 0.1982 $\pm$0.0091 0.3972 $\pm$0.1467 0.2131 $\pm$0.0386 0.3408 $\pm$0.1139 0.1825 $\pm$0.0435 0.0641 $\pm$0.0032 0.0524 $\pm$0.0022
GPF-plus 0.3109 $\pm$0.1086 0.2046 $\pm$0.0736 0.5239 $\pm$0.0720 0.3112 $\pm$0.0626 0.2112 $\pm$0.0102 0.1808 $\pm$0.0216 0.5042 $\pm$0.0559 0.3323 $\pm$0.0347 0.3606 $\pm$0.1570 0.2559 $\pm$0.0941 0.0687 $\pm$0.0078 0.0519 $\pm$0.0031
All-in-One 0.2337 $\pm$0.0878 0.0982 $\pm$0.0418 0.4310 $\pm$0.2792 0.2317 $\pm$0.1106 0.2116 $\pm$0.0217 0.1902 $\pm$0.0203 0.3127 $\pm$0.2199 0.1616 $\pm$0.1077 0.3831 $\pm$0.1751 0.1829 $\pm$0.0666 0.0559 $\pm$0.0251 0.0152 $\pm$0.0057
ProNoG 0.3366 $\pm$0.0349 0.2218 $\pm$0.0371 0.2789 $\pm$0.0466 0.2060 $\pm$0.0290 0.2154 $\pm$0.0083 0.2064 $\pm$0.0113 0.2620 $\pm$0.1002 0.2155 $\pm$0.0666 0.2986 $\pm$0.1067 0.1869 $\pm$0.0382 0.0748 $\pm$0.0159 0.0566 $\pm$0.0144
@@
E.4 Results of Baselines with Other Backbones
@@
In this section, we present the results of the baselines that use a GraphTransformer model UniMP Shi et al. [2020] and our used BWGNN Tang et al. [2022] as their backbone. To avoid excessive computational costs from grid searches, we utilize optimal pre-training and prompt-tuning combinations from Tables 9, Table 10, and Table 11. The results of transductive learning are shown in Table 12 and 14, and those of inductive learning are presented in Table 13 and 15. The results indicate that our model outperforms the baselines with various backbones.
E.5 Ablation Study